OpenStudy (anonymous):

What is the solution set for |3x – 24| less than or equal to 27?

OpenStudy (ash2326):

|3x-24|<=27 or -27<=3x-24<=27 -3<=3x<=51 -1<=x<=17

OpenStudy (anonymous):

you mean 16 right

OpenStudy (ash2326):

no, the solution set is x belongs to [-1,17]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Umm is this slope? just want to verification